### v2

# An Adaptive Language-Agnostic Pruning Method for Greener Language Models for Code

Abstract.
Language models of code have demonstrated remarkable performance across various software engineering and source code analysis tasks. However, their demanding computational resource requirements and consequential environmental footprint remain as significant challenges. This work introduces Alpine, an adaptive programming language-agnostic pruning technique designed to substantially reduce the computational overhead of these models. The proposed method offers a pluggable layer that can be integrated with all Transformer-based models. With Alpine, input sequences undergo adaptive compression throughout the pipeline, reaching a size that is up to less their initial size, resulting in significantly reduced computational load. Our experiments on two software engineering tasks, defect prediction and code clone detection across three language models CodeBert, GraphCodeBert and UniXCoder show that Alpine achieves up to a 50% reduction in FLOPs, a 58.1% decrease in memory footprint, and a 28.1% improvement in throughput on average. This led to a reduction in CO2 emissions by up to %. Importantly, it achieves a reduction in computation resources while maintaining up to 98.1% of the original predictive performance. These findings highlight the potential of Alpine in making language models of code more resource-efficient and accessible while preserving their performance, contributing to the overall sustainability of their adoption in software development. Also, it sheds light on redundant and noisy information in source code analysis corpora, as shown by the substantial sequence compression achieved by Alpine.
1. Introduction
Transformer-based language models (Vaswani et al., 2017) were initially conceived for natural language understanding. Under the naturalness hypothesis (Hindle et al., 2012), they were applied to solve diverse software engineering (se) tasks with remarkable accuracy. These tasks include code generation (Lai et al., 2023; Zeng et al., 2022a; Dibia et al., [n. d.]; Chen et al., 2022), quality assurance (Alqarni and Azim, 2022; Ciborowska and Damevski, 2023; Fatima et al., 2023; Tang et al., 2023), maintenance (Tian et al., 2023; Zhang et al., 2023; Le-Cong et al., 2023; Paul et al., 2023), and requirements engineering (Moharil and Sharma, 2023; Ezzini et al., [n. d.]).
Though the language models have shown exceptional promise, fine-tuning or running inference on these models require significant computational resources. Moreover, the demand of the resources is continuously increasing with the size of the language models. This entails that their training grows correspondingly, leading to increased usage of cloud computing services which operate on a pay-per-use basis. This inflates the direct financial investment for researchers and average users with consumer-grade hardware. In addition to the financial impact, the carbon footprint of training and maintaining these models is a growing environmental concern as they contribute to higher CO2 emissions (Luccioni et al., 2023; Lacoste et al., 2019; Wang et al., 2023b). These issues underscore the urgency for more resource efficient algorithms.
Shi et al. (Shi et al., 2023) have proposed Compressor as a step towards this direction. Compressor combines task-specific knowledge distillation (Hinton et al., 2015) and evolutionary search of network architecture to generate a configuration that yields a smaller model. Task-specific knowledge distillation often falls short in terms of task-specificity and computational overhead. The distilled student models are typically optimized for a specific task or domain, limiting their generalization abilities. Moreover, the process of task-specific knowledge distillation itself can be computationally expensive, as it involves fine-tuning the teacher model either way, generating predictions or soft labels, pre-training, and fine-tuning the student model.
Similarly, DietCode (Zhang et al., 2022) attempts to achieve better resource utilization by assigning importance scores to tokens and statements based on attention weights and then selecting the most important statements to form a new, condensed input. However, it has two main limitations. First, it relies on the peculiarities of the task’s programming language, such as specific syntax, semantics, and constructs. This reliance makes extending the method to all programming languages difficult without overhead and customization. Second, it assumes that the attention weights remain static and applicable across different corpora and tasks. However, this is not the case given that these attention weights are derived from trainable parameters, which are updated when the language model is fine-tuned on other tasks.
LTP (Kim et al., 2022) covers some of DietCode’s limitations. It proposes a method that learns a threshold at each Transformer layer and tokens are pruned if their importance score is below such threshold. Concretely, the pruning is done by zeroing out the hidden representations of these tokens. However, such an approach does not effectively reduce computation. While the method assigns zero values to pruned tokens by multiplying the hidden representations with a mask111LTP’s implementation: https://t.ly/qpu4J, the sequence length remains unchanged. Consequently, these zero-valued representations still participate in subsequent matrix multiplications and other operations. In most hardware and software implementations, multiplications involving zero are not automatically optimized out. As a result, the same number of floating-point operations (FLOPs) are performed regardless of how many tokens are “pruned” leading to no actual reduction in computational cost. Finally, LTP requires multiple fine-tuning steps: in the first pass, it learns the thresholds as well as full model parameters, then fine-tunes the model’s weights in the second pass using those thresholds.
To tackle the aforementioned limitations, we present Alpine, a language-agnostic token pruning technique seamlessly integrated into any Transformer-based model. This approach involves placing a pruning module within the Transformer block. Alpine processes a code snippet, generating a set of importance scores for each token. Based on these scores and a configurable importance range, it completely eliminates tokens that exist outside the range. Consequently, the reduced input size yields considerable speed up in both fine-tuning and inference phases. Notably, Alpine operates without requiring supplementary code information, such as statement types, ensuring its programming-language independence. Furthermore, the importance scores are computed during the fine-tuning process, enabling it to adapt dynamically to various downstream Software Engineering (SE) tasks.
We evaluate our proposed method on two software engineering tasks defect prediction and code clone detection using three language models, CodeBert, GraphCodeBert and UniXCoder. Our approach allows these models to use less computation, measured by the number of floating point operations (flops) and memory footprint, with minimal loss in performance and even slight improvement in some scenarios.
This study makes the following contributions to the field.
We conduct an analysis to quantify the computational requirements of the individual components within the Transformer model measured by flops. The analysis helps us identify the bottleneck components and understand the changes in computational complexity with respect to the input sequence length.
Using these insights, we design an effective pruning method that makes language models computationally efficient. Its plug-and-play characteristic makes it easy to integrate and use with virtually any Transformer-based model with minimal overhead.
We illustrate how Alpine can maintain a considerable amount of the original accuracy exhibited by the none-pruned models with substantially less computation and carbon emission.
Alpine’s ability to reduce computational requirements while maintaining accuracy has two significant implications.
First,
it contributes to more sustainable software development practices by reducing the required computation and carbon footprint associated with such models.
Second, it lowers the barriers to entry for developers and researchers, facilitating the adoption of language models for code, especially when constrained by consumer-grade gpus. As a result, it offers a more accessible and environmentally friendly solution for the software engineering community.
Replication package: Our replication package including source code and data is available online (Saad et al., [n. d.]).
2. Background and Motivation
In this section, we provide a brief primer on the Transformer architecture (Vaswani et al., 2017) and the implications of its computational requirements. Then, we investigate the number of operations performed by the main blocks of this architecture to identify a potential bottleneck within a certain class of Transformer-based models.
2.1. Transformer Architecture
The Transformer architecture, introduced by Vaswani et al. (Vaswani et al., 2017) for natural language processing (nlp), is a sequence-to-sequence model composed of two main components: an encoder and a decoder. This work focuses on the encoder-only variant, one of the most used variants in SE-related tasks (Hou et al., 2023b). We further elaborate in Section 6.3 how Alpine can be extended to encoder-decoder and decoder-only models. For the rest of the paper, we refer to a Transformer encoder layer by a Transformer layer for convenience.
Input Representation: The input fed into the Transformer layer consists of the tokenized sequence of tokens and an attention mask . includes a special [CLS] token, whose representation is used for classification or ranking tasks, and a [SEP] to indicate the end of the sequence, followed by padding tokens [PAD], if necessary222Appending two sequences is also another format that can be used as input: I.
The attention mask is used to indicate which tokens should be attended to and which ones should be ignored,
The input representation is then passed through an embedding layer to obtain dense vector representations for each token, which serve as the input to the encoder layer.
Main blocks: The main blocks of this layer are the Multi-head Self-Attention (mha) which is composed of attention heads, and the Position-wise Feed-forward network (ffnn). mha is defined as follows
where is the scaled dot attention function introduced by Vaswani et al. (Vaswani et al., 2017), and is the input sequence. The derivation of Equation 1 is taken from the work of Kobayashi et al. (Kobayashi et al., 2020) and Saad and Sharma (Saad and Sharma, 2023). After the mha, the ffnn block is applied to each position of the sequence independently. It is a subnetwork composed of two linear layers with the application of a non-linear activation function ReLU in between.
Note that, at the end of each block, the Transformer layer employs residual connections followed by a layer normalization to facilitate the flow of information and stabilize the training process.
2.2. Motivation
2.2.1. Impact on Energy Consumption
Language models for code require a significant amount of computational power to train and run. The number of parameters in these models has increased from millions to billions (Hou et al., 2023a; Zhao et al., 2023), necessitating the use of high-performance computing systems and specialized hardware, such as graphics processing units (gpus) and tensor processing units (tpus). The training of such models involves processing large amounts of data, often in the order of terabytes. This requires not only powerful processors but also high-bandwidth memory and storage systems. As a result, the energy consumed during the training phase is substantial, as it involves running these models for extended periods that can reach the magnitude of days. Table LABEL:tab:model_training_time illustrates the training time of different models of various sizes.
| Model | # Parameters | Training Time |
| UniXCoder (Guo et al., 2022) | 125M | 8 days |
| CodeT5base (Wang et al., 2021) | 220M | 12 days |
| InCoder (Fried et al., 2023) | 6.7B | 24 days |
The energy consumption of language models is directly proportional to their computational requirements. As the models become more complex and the amount of data they process increases, the energy needed to power the computing systems also rises. This results in a higher carbon footprint of these models. Wang et al. (Wang et al., 2023b) have shown that the carbon footprint of pre-training the bert model (Devlin et al., 2018) on a machine with one NVIDIA Quadro RTX gpu is kg CO2. This is equivalent to the same amount emitted by driving a vehicle for a distance of kilometers333Wang et al. (Wang et al., 2023b) report the distance in miles (Wang et al., 2023b). Moreover, we are witnessing a notable trend of models increasing in size leading to even higher environmental impact. For instance, recently, Lozhkov et al. (Lozhkov et al., 2024) released StarCoder 2, a family of language models for code with sizes ranging from B to B parameters. They reported that the smallest model has resulted in kg of CO2 emission during pre-training.
2.2.2. Computational Cost of the Transformer Layer
This section explores the computational complexity and memory requirements of both the mha and ffnn layers.
The computational cost of a neural layer can be measured in terms of the number of floating-point operations (flops) performed by such layer. We calculate the theoretical operations count, i.e., the upper bound when a sequence has the length of the maximum number of tokens. The flops counting rules introduced by Hunger (Hunger, 2005) for matrix multiplications are used in our work. Specifically, the rule stating that the number of flops of a matrix multiplied with another matrix is . Note that here and represent the dimensions of the matrices and , respectively.
FLOPs of the mha: Let be the maximum number of tokens, the hidden dimension, and the number of heads. The first step includes three linear projections of the input with the , , and weight matrices at each attention head. These operations involve matrix multiplication and addition (for the bias term). The total number of operations at this stage is calculated as follows.
The scaled dot-product attention is calculated by multiplying the resulting query matrix with the transpose key matrix coupled with the application of the scaled softmax function.
The next operation in the mha layer is the multiplication of the attention probability matrices of each head obtained in Equation 8 with the value matrix that was calculated from one of the linear projections performed earlier.
Finally, the outputs of all attention heads are concatenated and projected through a linear layer. The concatenation operation requires 0 operations, hence, the only flops performed are from the matrix multiplication of the concatenated matrix with .
Summing the Equations 4-10, we get the following formula to determine the flops count of the attention layer.
FLOPs of the ffnn: The ffnn takes the output of the mha’s layer as input, scales it to a higher dimension than the mha’s, applies a gelu transformation (Hendrycks and Gimpel, 2016)444While the Transformer layer was stated earlier to use ReLU activation, the models used in this work replace it with the gelu function. Note that this does not affect the calculations performed above., and scales it back to the original hidden dimension.
Following the same notation, let be the number of tokens in a sequence and the dimension of the ffnn layer. The flops count for the first linear layer includes a multiplication between the output matrix of the mha with the weight matrix of this layer proceeded by the application of the gelu function.
The final operation projects the output to the original hidden dimension.
The ffnn’s dimension is usually set to be higher than the mha’s. For instance, in the models that we have chosen for our experiments, . Using this fact, and the Equations 12 and 13 the total number of flops is,
Although Equations 11 and 14 show that the computational cost of mha increases quadratically compared to the ffnn, it uses fewer flops when the maximum sequence length is . Such length is an upper bound for many language models such as CodeBert (Feng et al., 2020), GraphCodeBert (Guo et al., 2021), and UniXCoder (Guo et al., 2022) that are used to solve SE tasks. We now prove this statement analytically,
Lemma 0.
For every conceivable input sequence fed into a Transformer model with hyperparameters same as CodeBert, GraphCodeBert, and UniXCoder, it holds true that
Proof.
Let us fix , and consider the following function:
This function
is quadratic
has a maximum .
has two roots and
therefore or when . In other words, if the sequence length , then
For UniXCoder, GraphCodeBert, and CodeBert, we have that and . Thus,
Furthermore, the maximum sequence lengths are 1024, 512, and 512, for UniXCoder, GraphCodeBert, and CodeBert respectively. Therefore, it will consistently hold true that , leading to the conclusion that .
This is further demonstrated empirically when measuring the number of flops performed by both of these layers in CodeBert (Feng et al., 2020) using the test set of the Devign (Zhou et al., 2019) dataset as shown in Figure 1.
Therefore, reducing the sequence length before feeding it to the ffnn layer is a potential way to reduce such complexity.
3. Approach
In this section, we present the pruning procedure adopted in this study in each Transformer layer and elaborate on the design choices.
As shown in Figure 2 and Algorithm 2, the pruning layer is inserted between the mha and ffnn layers. As highlighted in Lemma 2.1, the feed-forward layer represents a bottleneck in terms of computational costs and memory consumption (Ganesh et al., 2021) caused by the projection of the attention layer’s output to a higher dimension in the ffnn. Hence, performing pruning before this layer would reduce the effect of such a bottleneck.
Algorithm 1 presents the steps to perform the pruning. The first step is to assign a score to each token to determine its importance within the sequence. In this work, we use attention based scoring (Kim et al., 2022), which quantifies this importance using the attention probabilities output by the mha layer. Specifically, in Lines 4–11 we take the mean across the attention matrices of the attention heads of the mha layer and then perform a column-wise mean to get a distribution of importance scores as illustrated by Equation 15, where is the number of attention heads, is the attention matrix of an attention head and is the list of tokens excluding the special tokens [CLS], [SEP], and [PAD].
The reason we exclude the [CLS] and [SEP] tokens from pruning is two-fold. First, the [CLS] token representation is extracted from the output of the final transformer encoder layer and fed into a linear classification layer as done previously (Shi et al., 2023; Lu et al., 2021). Hence, such a token should be kept throughout the encoder pipeline. Second, it has been shown that these two tokens usually receive the highest attention scores (Clark et al., 2019; Kobayashi et al., 2020; Saad and Sharma, 2023) in natural and programming languages, which would skew the importance scores distribution. As for the padding token, its attention score is 0, which again would affect the score distribution, especially for shorter sequences.
Once we obtain the importance scores distribution, we keep the tokens that are within a specific range. We set the range to be , where and are the mean and standard deviation of the importance scores distribution, and is a hyperparameter that defines the width of the lower and upper bounds as shown in Lines 12-16. The smaller gets, the tighter the window, the less tokens are kept. This hyperparameter controls the aggressiveness of pruning.
Based on a preliminary investigation, we observed that the distribution of importance tokens follows roughly a leptokurtic distribution (i.e., distribution with low standard deviation and a high peak) at each layer. By removing tokens outside the defined range in such distribution, we can preserve the most informative and representative data while reducing computational complexity.
Given the importance scores and the range , we create a new mask that indicates the tokens that should be pruned and those that should be kept (Lines 17-20),
Algorithm 2 shows the modified implementation of the Transformer layer. The highlighted Lines 7 and 8 represent the modifications that we have introduced. This also demonstrates the ease of integrating Alpine with minimal effort.
Using the updated mask from the previous step, we either remove the rows from the mha output matrix that correspond to the tokens to be pruned as shown in Line 8, or merge them into one row. By default, we do the merging to minimize the information loss incurred by pruning, especially that of tokens that are above the upper bound of the importance interval. Next, the matrix’s dimensions are reduced. We also calculate the final mask since this operation is performed on a batch of sequences to guarantee that all sequences are of equal length.
We experiment with three variants of pruning. The first one involves performing pruning across all layers of the encoder. In the second setting, it is performed only for even-indexed layers (i.e., ). Whereas the final one involves pruning at odd-indexed layers (i.e., ).
4. Experiments
The goal of this study is to investigate the effect of Alpine on language models of code when applied to SE tasks. The research aims to assess the balance between efficiency, from computational and environmental points of view, and model accuracy. Toward this goal, we formulate the following research questions. The rest of the section elaborates on the experiments designed to answer the research questions.
RQ1.
To what extent does pruning on language models for code save computational resources?
Intuitively, the less data a model processes, the more computation is saved. With this question, we wish to confirm and quantify these computational savings compared to the baseline, non-pruned model. -
RQ2.
What is the impact of the pruning technique on the performance of language models for code on various SE tasks?
Naturally, pruning tokens would result in some information loss. In this research question, we aim to measure the extent of performance drop, in terms of accuracy, for example, that might occur, if any. -
RQ3.
What is the effect of merging prunable tokens as opposed to completely dropping them?
The proposed approach allows one to choose whether pruned tokens are to be partially kept by merging their representation into a single row or entirely removed from the sequence. In this research question, we investigate the impact of such a design choice. -
RQ4.
What is the impact of the computing environment on the efficiency of the pruning method?
Finally, an important reason behind model simplification is to allow computationally demanding models to run on relatively lesser-capable gpus (e.g., consumer-grade gpus) compared to those used in high-performance computing clusters (e.g., NVIDIA V or A). This exploration would ensure wider accessibility and make using these models more practical.
4.1. Software Engineering Tasks
In this study, we choose two substantially different tasks from the CodeXGLUE benchmark (Lu et al., 2021) for our experiments. Both tasks aim to improve code quality, developer productivity, software reliability, and maintainability.
Code clone detection: Clone detection identifies similar or identical code fragments, helping maintain code quality, reduce bugs, and enable refactoring. In this work, we used the filtered version of the BigCloneBenchmark dataset (Svajlenko et al., 2014) released by Wang et al. (Wang et al., 2020). This binary classification problem aims to predict whether two Java code snippets are clones. The dataset is composed of training pairs and pairs for testing. We randomly take a subsample of pairs for training and for evaluation; such random sampling is quite common in this domain (Shi et al., 2023; Yang et al., 2022). We use the same subsamples for all experiments during training and evaluation.
Defect prediction: It identifies bugs, vulnerabilities, and defects early, reducing the cost and effort of manual code reviews and testing. We used Zhou et al. (Zhou et al., 2019)’s dataset that was curated from a set of four large-scale open-source C repositories. The authors manually annotated the code snippets and covered multiple types of defects such as resource leaks and use-after-free vulnerabilities. We use the default pre-split sets of samples for training and for validation and testing.
4.2. Model selection
We select the following models given that they represent state-of-the-art encoder-only models in solving the previously mentioned tasks as of conducting this work (Microsoft Research, 2023; Shi et al., 2024; Niu et al., 2023; Zeng et al., 2022b).
CodeBERT (Feng et al., 2020): It is a pre-trained language model that has the same architecture and a similar pretraining strategy as RoBERTa (Liu et al., 2019). It comprises twelve Transformer layers, each containing a 12-headed mha layer and has a hidden dimension of . It was trained on bimodal data consisting of pairs of natural language and programming language across six programming languages from the CodeSearchNet dataset (Husain et al., 2019).
GraphCodeBERT (Guo et al., 2021): GraphCodeBert extends CodeBert by including the data flow information within the source code input. It also includes an additional pre-training task where the objective is to predict whether a data flow edge exists between two nodes.
UniXCoder (Guo et al., 2022): UniXCoder leverages a multi-modal input consisting of natural language, source code, and the flattened sequence of the abstract syntax tree of the code snippet. It is similar to the two other models that share the same internal architecture. The difference here (aside from the pre-training objectives and input representation during pre-training) is that UniXCoder allows for a larger context length, compared to for (Graph)CodeBERT.
We use the same model architecture in both tasks. Specifically, we add a classifier in the form of a dense layer on top of each encoder. It takes as input the last hidden representation of the [CLS] token.
4.3. Evaluation Metrics
To assess the predictive performance of the models on the aforementioned tasks, we use the same evaluation metrics that were reported by Lu et al. (Lu et al., 2021) in the CodeXGLUE benchmark. Specifically, we calculate the accuracy for the Devign dataset and F1-score for the BigCloneBenchmark. As for the computational efficiency, we report the number of floating points operations (flops) and Throughput. As stated in Section 2.2 the number flops quantifies the computational complexity. A model with a higher flop count requires more computing resources. The goal is to reduce such a count while maintaining the predictive performance as much as possible. The throughput refers to the number of input samples a model can process in a second. This metric is especially relevant during inference and model deployment.
4.4. Experimental Setting
We ran the experiments on two machines. The first machine has an amd Milan 7413 cpu, gb of ram and an NVIDIA A with gb of vram. The second machine has an Intel i cpu, gb of ram, and an NVIDIA RTX gpu with gb of vram. We use the same hyperparameters set in the CodeXGLUE benchmark for each task.
| CodeBert | GraphCodeBert | UniXCoder | |||||||
| Pruning Method | Accuracy | FLOPs () | TP | Accuracy | FLOPs () | TP | Accuracy | FLOPs () | TP |
| Baseline (No Pruning) | % | % | % | ||||||
[…]
| Even Indexed | % | % | % | ||||||
| Odd Indexed | % | % | % | ||||||
| CodeBert | GraphCodeBert | UniXCoder | |||||||
| Pruning Method | F1 | FLOPS () | TP | F1 | FLOPS () | TP | F1 | FLOPS () | TP |
| Baseline (No Pruning) | % | % | % | ||||||
[…]
| Even Indexed | % | % | % | ||||||
| Odd Indexed | % | % | % | ||||||
5. Results
5.1. RQ1: Computational Efficiency
Impact on FLOPs Count: Table 2(b) highlights the computational speedup offered by Alpine across the three models on the SE tasks. Performing pruning across all layers allows for the most computational savings. On average555Given the values are ratios, we report the geometric mean., it reduces the flops666Calculated using fvcore. count by on the Devign dataset and and BigCloneBenchmark compared the baseline models (i.e., without any pruning). Specifically, CodeBert and GraphCodeBert exhibit roughly the same gain, where these models consume roughly half flops. UniXCoder also show significant gain, though slightly smaller ( and ), in terms of flops.
Interestingly, performing pruning at even indexed layers seems to yield lower operations compared to pruning at odd indexed layers. A possible explanation for this behavior is that there is a considerable drop in the number of tokens at the 0th layer which would have an earlier cascading effect if we were to perform this at a later layer.
In Figure 3, we plot the average sequence length across all the twelve layers during inference. When pruning across all layers on the Devign dataset, the sequence length plot exhibits an exponential-like decay on CodeBert and GraphCodeBert. On the other hand, this reduction follows roughly a linear trend on BigCloneBenchmark across all models. What is interesting is that at the final layer, the sequences are quite compressed. Specifically, the highest average sequence length at the last layer () is obtained when using GraphCodeBert on the defect prediction task which is less than the none pruned average length of . In addition, aligned with the results of Table 2(b), we can see that pruning at even-indexed layers yields shorter sequences than odd-indexed ones. Across all task-model combinations and all layers (except for - layer), Alpine produces more compressed sequences when applied at these layers compared to their odd-indexed counterpart.
Impact on Throughput: Pruning across all layers consistently yields the highest throughput improvement for all models and tasks. For instance, CodeBert’s throughput increases from to (% improvement) for the defect prediction task and from to (% improvement) for code clone detection. Moreover, applying pruning in even-indexed layers generally results in better throughput compared to pruning odd-indexed layers. This trend is evident across all models and tasks, with even-indexed layer pruning providing an average throughput improvement of % compared to no pruning, while odd-indexed layer pruning offers an average improvement of %.
Another observation is that the extent of improvement varies among the models. GraphCodeBert exhibits the highest average throughput improvement of % across all pruning methods and tasks, followed by CodeBert with % and UniXCoder with %. In general, the models achieved higher throughput on the Devign dataset, regardless of the pruning method applied. The average throughput for this task is samples per second, while for the code clone detection task, it is samples per second.
The increase in throughput is attributed to Alpine’s ability to compress sequences. A smaller sequence length allows for more efficient parallel processing and reduces the computational overhead associated with longer sequences, resulting in higher throughput and faster processing times.
Impact on Memory Footprint: To further evaluate Alpine’s impact from a computational efficiency standpoint, we report the gpu memory consumption. Due to space constraints, we only report the figures regarding pruning across all layers.
From the plot in Figure 4, we see that Alpine significantly reduces memory consumption across all models and tasks. CodeBert exhibits the highest percentage of memory reduction for the defect detection task (%), while GraphCodeBert and UniXCoder show more consistent reductions across both tasks (% to %).
This reduction leads to improved cost efficiency, faster data transfer, and power efficiency. By requiring less memory, models can be deployed on more affordable gpus with smaller memory capacities, reducing hardware costs. Additionally, a smaller memory footprint results in faster data transfer between the cpu and gpu; the additional speed, in turn, increases throughput as demonstrated earlier. Moreover, minimizing the model’s memory footprint contributes to reduced power consumption, which is crucial in constrained environments.
5.2. RQ2: Impact on Performance
In Table 2(b), we report the predictive performance of the pruned models. On the Devign dataset, performing pruning at odd-indexed layers leads to a slight improvement when using CodeBert, with a % increase in accuracy compared to that of the non-pruned version. On the other hand, adding the pruning layer at even layers is more efficient for GraphCodeBert and UniXCoder, where they were able to maintain % and % of the original accuracy scores, respectively. In the clone detection task, pruning at even-indexed layers achieves the highest F1-scores compared to other pruning strategies. It even managed to improve the clone detection performance by increasing the F1-scores of GraphCodeBert and UniXCoder by % and %.
Applying Alpine at all layers not only leads to less computation but also results in a high level of predictive performance for all three models. UniXCoder exhibits the highest performance both before and after pruning, retaining % of its original accuracy. CodeBert and GraphCodeBert also demonstrate strong performance retention, preserving % and % of their original accuracy, respectively. A similar observation can also be made for the results of code clone detection. All models maintain a high level of performance after pruning, with UniXCoder achieving the highest F1-score retention at % of its original score. GraphCodeBert and CodeBert also demonstrate strong performance, retaining % and % of their original F1-scores, respectively.
Among the three models, UniXCoder shows the highest resilience to pruning, keeping an average of % of its original performance across both tasks and all pruning methods.
5.3. RQ3: Effect of Merging Prunable Tokens on Model Accuracy
| Model | Task | |||||||||||
| Defect Prediction | Code Clone Detection | |||||||||||
| All | Even | Odd | All | Even | Odd | |||||||
| w/ m | w/o m | w/ m | w/o m | w/ m | w/o m | w/ m | w/o m | w/ m | w/o m | w/ m | w/o m | |
| CodeBert | % | % | % | % | % | % | % | % | % | % | % | % |
| GraphCodeBert | % | % | % | % | % | % | % | % | % | % | % | % |
| UniXCoder | % | % | % | % | % | % | % | % | % | % | % | % |
During pruning, the representation of tokens marked to be pruned can be entirely removed or merged into one vector. To study the impact of token merging, we conduct experiments to measure its impact on the two tasks and report the results in Table 3.
Across all models and datasets, merging tokens consistently yields better performance compared to discarding them completely. UniXCoder benefits the most from this design choice, with an average performance improvement of % on Devign and % on BigCloneBenchmark.
On the Devign dataset, GraphCodeBert shows the highest performance gain when pruning even layers (%), while CodeBert’s performance is enhanced the most when pruning odd layers on the BCB dataset (%). In contrast, GraphCodeBert exhibits the least improvement when pruning all layers on the BigCloneBenchmark dataset (%).
5.4. RQ4: Impact of Computing Environment
In Figures 6 and 6, we plot the time taken to fine-tune the language models on the selected tasks when using different gpus. First, examining the results for the A gpu, we observe that Alpine leads to a reduction in fine-tuning time for all three models on both the tasks. For defect prediction, pruning results in a running time reduction of around % for CodeBert, % for GraphCodeBert, and 15% for UniXCoder. Similarly, for the clone detection task, pruning reduces the running time by approximately % for CodeBert, % for GraphCodeBert, and % for UniXCoder. These findings demonstrate that pruning is effective in improving the efficiency of the models on a high-performance gpu.
Next, comparing the results between the RTX and A gpus, we notice that the running times on the A are generally shorter than those on the RTX, both with and without pruning. This is expected due to the superior computational capabilities of the A. However, the relative impact of pruning on efficiency differs between the two gpus. On the RTX, pruning leads to a more significant reduction in running time, with improvements ranging from % to % across models and tasks. In contrast, on the A, the reduction in running time due to pruning is less pronounced, ranging from % to %.
| gpu | E(CO2 kg) without Alpine | E(CO2 kg) with Alpine | Emission reduction rate |
| RTX | % | ||
| A | % |
Building upon these results, we further investigate the implications of the reduced time from an environmental perspective.
In Table 4, we report the CO2 emission rates for both gpus with and without Alpine. The measurements were taken using the tool provided by Lacoste et al. (Lacoste et al., 2019).
The results show that Alpine significantly reduces the CO2 emissions of both the RTX and A gpus. For the RTX, the amount of CO2 emitted decreases from kg to kg with Alpine, resulting in a substantial emission reduction rate of %. Similarly, for the A, emissions decrease from kg to kg, resulting in a reduction rate of %. The reduction in emissions is particularly significant for the RTX, which aligns with the previous analysis showing that pruning leads to a more substantial improvement in efficiency on consumer-grade gpus compared to high-performance gpus like the A.
6. Discussion
6.1. Sequence Compression
In Section 5.1 we demonstrated how Alpine can significantly reduce sequences’ lengths to less than the original lengths in some scenarios. In code analysis tasks, not all parts of the input sequence contribute equally to the final prediction. Many tokens in source code, such as common language keywords, special characters, or even standard library imports, may not be critical for identifying vulnerabilities or detecting code clones. However, as previously stated in Section 5.1, the reduction rate depends on the nature of the task.
Vulnerability detection, which is the focus of the Devign dataset, usually involves identifying specific code patterns that may lead to security issues. These vulnerability indicators often compromise a small portion of overall function, allowing for rapid pruning of irrelevant code segments. The exponential-like decay observed in Figure 3 suggests that the models may quickly isolate potentially vulnerable code portions, discarding a large amount of non-critical code in earlier layers.
On the other hand, in code clone detection, the task associated with BigCloneBenchmark, requires a more holistic view of code structure and functionality. Clones can manifest in various forms, from exact verbatim copies (i.e., Type-1 clones) to functionality-similar but syntactically different implementations, also known as Type-4 clones. This entails the need to retain a larger set of code features through the layers, which explains the more gradual, linear-like reduction in sequence length. The models likely need to maintain a wider context to effectively compare and identify similarities across different code snippets.
6.2. Less is More
Our findings on sequence reduction with Alpine caters to the calls of developing smaller, task-oriented models (Menzies, 2023). Just as compact models demonstrate that not all parameters in large, general-purpose models are necessary for specific tasks, our results show that not all tokens in an input sequence are crucial for effective code analysis tasks. In both cases, the underlying principle is the same: task-specific efficiency. (Large) language models, while powerful, often contain redundant or task-irrelevant information. Similarly, full code sequences include many tokens that may not contribute significantly to solving SE tasks. This is also similar to previous research. For instance, Xu et al. (Xu et al., 2023b) have observed that removing noisy instances when training models for obsolete comment detection can lead to up to % improvement in accuracy, with % less data. The findings of this work align with the principle of less is more. We can achieve efficient, high-performing systems not by processing more data, but by intelligently selecting the most pertinent information and avoiding noisy data for the task at hand.
6.3. Extension of Alpine to other Transformer Variants
The main focus of Alpine was encoder-only models given their ability to generate embeddings that are primarily suitable for identification tasks. They are also an important component in text generation systems that use information retrieval, or Retrieval Augmented Generation (RAG). However, it is important to remark that Alpine can be extended to other variants of the Transformer architecture, i.e., decoder-only and encoder-decoder models. In decoder-only models, Alpine would be placed after the Masked Self-attention layer (MaskedSA), not after the self-attention layer because the MaskedSA attends only to previous tokens, preserving the autoregressive nature of the decoder. In addition, the last token would be kept, as currently done, so that it can be used by LM head. Finally, in the encoder-decoder setting, Alpine can be either applied to the encoder, decoder, or both. The evaluation of Alpine’s performance on these architectures is deferred to future work. In return, this would increase their adoption given the less computational resources they demand, especially on consumer-grade hardware.
7. Threats to Validity
Internal validity:
Internal threats to validity are concerned with the ability to draw conclusions from our experimental results.
The hyperparameters specified for a model influence the time, resources required, and the results.
To mitigate a potential threat, we used the same hyperparameters across all models, as reported in previous works (Lu et al., 2021), and utilized the same data splits for fine-tuning and testing.
This consistency in the experimental setup strengthens the internal validity of our findings, ensuring that the observed effects can be attributed to the Alpine pruning method rather than variations in hyperparameters or data splits.
External validity:
Such threats are concerned with the ability to generalize our results.
The language-agnostic nature of the Alpine pruning method allows it to be applied to a wide range of programming languages without the need for language-specific adaptations.
We evaluate the approach on three language models and tasks with different programming languages (C and Java).
Its compatibility with all transformer-based encoder models makes it trivial to integrate into various architectures, enhancing its generalization.
Finally, the two gpus used in this study, NVIDIA A and RTX, belong to Ampere and Turning architecture families. These architectures encompass a larger set of other cards which we expect to exhibit the similar reported trends.
Construct validity:
Construct threats to validity are concerned with the degree to which our analyses measure what we claim to analyze.
We employed well-established metrics for measuring predictive performance (such as, accuracy and F1-score),
computational efficiency (flops, throughput, running time and memory footprint),
and environmental impact through CO2 emissions rate.
Relying on the traditional metrics for the measured aspects mitigates the potential threats of construct validity and ensures that the variables are accurately captured.
8. Related Work
8.1. Pre-trained language models of code
Pre-trained code language models have demonstrated benefits in various SE tasks, including code search, code and documentation generation, and defect prediction (Wang et al., 2021; Guo et al., 2022; Lu et al., 2021; Sharma et al., 2024). Typically based on the transformer architecture (Vaswani et al., 2017), these models can broadly be categorized into decoder-only, encoder-only, and encoder-decoder language models (Min et al., 2023).
Decoder-only architectures are typically pre-trained in an autoregressive manner, i.e., predicting the subsequent token based on preceding ones. In the realm of code, examples include InCoder (Fried et al., 2023), StarCoder 2 (Lozhkov et al., 2024), and CodeGen 2 (Nijkamp et al., 2023). These architectures excel in code generation tasks.
Encoder-only architectures are primarily pre-trained on the masked language modeling task i.e., these models are trained to predict masked words given the remaining sequence as context. Examples include CodeBert (Feng et al., 2020) and GraphCodeBert (Guo et al., 2021). They are frequently used for classification tasks or as a means of feature extraction.
Encoder-decoder architectures undergo pre-training through conditional generation tasks i.e., generating text based on provided input text, e.g., for code summarization or generating code from natural language descriptions. Sample architectures include CodeT5 (Wang et al., 2021), CodeT5+(Wang et al., 2023a), and plbart (Ahmad et al., 2021).
8.2. Optimized transformers
The techniques to enhance the effectiveness of a transformer model while maintaining its performance fall into three categories—knowledge distillation, quantization, and pruning (Shi et al., 2023; Kim et al., 2022; Zhu et al., 2023).
Knowledge distillation. This technique trains a smaller model (the student model) to imitate the behavior of a larger model (the teacher model) (Gou et al., 2021). In a SE context, Shi et al. (Shi et al., 2023) introduced Compressor, a framework that employs task-specific knowledge distillation to improve the efficiency of the final transformer model. The authors perform a neural architectural search via a genetic algorithm, followed by training the selected model using knowledge distillation. In their experiments, they use CodeBert (Feng et al., 2020) and GraphCodeBert (Guo et al., 2021) as teacher models to tackle the tasks of vulnerability prediction and code clone detection. Compressor still requires the fine-tuning of the teacher model, which incurs overhead throughout the process. Alpine allows to reduce the computational costs of fine-tuning these models, minimizing such an overhead.
Quantization. It converts the model’s parameters from 32-bit floating-point numbers (float32 data type) to lower-precision formats, such as 8-bit integers (int8 data type) (Nagel et al., 2021). Depending on the component of quantization, these techniques can be categorized as quantization of weights only (Xu et al., 2023a; Dettmers et al., 2024; Kim et al., 2023) and quantization of both weights and activations (Zafrir et al., 2019; Yao et al., 2022; Dettmers et al., 2023). Although quantization reduces the memory footprint of these models, it may also reduce the inference speed and accuracy (Jin et al., 2024).
Pruning. Pruning techniques can be divided into weight pruning and token pruning (Kim et al., 2022). Liu et al.(Liu et al., 2021) adopt weight pruning to prune unimportant heads and channels from the mha and ffnn respectively. Zang et al. (Zhang et al., 2022) offer DietCode—a token pruning approach. To build DietCode, Zang et al. carry out an empirical investigation to determine the significance of statements and tokens for CodeBert. Subsequently, based on these findings, they employ a 0-1 Knapsack approach to prune the input sequence. In contrast, our approach is performed adaptively at each layer. Additionally, during the pruning step, we could aggregate the pruned tokens into a single one to prevent excessive loss of information. Moreover, our method is language-agnostic, facilitating easy adaptation to diverse programming languages and extending applicability to natural language inputs, particularly beneficial in minimizing computational overhead in tasks like code search or code generation from documentation.
As for LTP (Kim et al., 2022), in addition to the main limitations previously mentioned where concrete computational speed-up is not achieved, there exist some limitations concerning the methodology. First, the reported number of FLOPs is manually calculated, under the assumption that matrix multiplication is optimized when zeroed-out vectors are present. Furthermore, given the placement of the pruning layer, these representations do not remain zeroed-out given the presence of skip connections and normalization layers777This was also pointed out by the issues raised in the replication package repository: https://t.ly/zBMGE and https://t.ly/kLCiU.. Moreover, it does not mask tokens during the evaluation,i.e., the full sequence, threatening the validity of the accuracy results.
9. Conclusions and Future Work
This study addresses the challenges, specifically computational efficiency and environmental impact, associated with language models of code. We propose Alpine, a pruning method that reduces the input sequence while maintaining predictive performance. Our experiments on three language models and datasets with different programming languages demonstrate that it significantly improves computational efficiency that, in turn, reduces CO2 emissions. The results show that it is particularly effective on consumer-grade gpus, enabling the usage of these models on more accessible hardware. Furthermore, the programming language-agnostic nature of Alpine and its compatibility with Transformer-based models enhance the generalizability of our findings. We envision to extend this work along two axes. First, to assess Alpine on encoder-decoder, and decoder-only models. Second, we aim to investigate other scoring importance functions and more adaptable merging strategies.
10. Data Availability
Our replication package including source code and data is available online through this repository (Saad et al., [n. d.]).
Acknowledgements.
This work is partially funded by by the Natural Sciences and Engineering Research Council of Canada (NSERC) through grant NSERC Discovery RGPIN/.References
- (1)
- Ahmad et al. (2021) Wasi Uddin Ahmad, Saikat Chakraborty, Baishakhi Ray, and Kai-Wei Chang. 2021. Unified pre-training for program understanding and generation. arXiv preprint arXiv:2103.06333 (2021).
- Alqarni and Azim (2022) Mansour Alqarni and Akramul Azim. 2022. Low Level Source Code Vulnerability Detection Using Advanced BERT Language Model. Proceedings of the Canadian Conference on Artificial Intelligence (may 27 2022).
- Chen et al. (2022) Fuxiang Chen, Fatemeh H. Fard, David Lo, and Timofey Bryksin. 2022. On the transferability of pre-trained language models for low-resource programming languages. In Proceedings of the 30th IEEE/ACM International Conference on Program Comprehension (ICPC ’22). 401–412. https://doi.org/10.1145/3524610.3527917
- Ciborowska and Damevski (2023) Agnieszka Ciborowska and Kostadin Damevski. 2023. Too Few Bug Reports? Exploring Data Augmentation for Improved Changeset-based Bug Localization. arXiv preprint arXiv:2305.16430 (2023).
- Clark et al. (2019) Kevin Clark, Urvashi Khandelwal, Omer Levy, and Christopher D Manning. 2019. What does bert look at? an analysis of bert’s attention. (2019).
- Dettmers et al. (2024) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2024. Qlora: Efficient finetuning of quantized llms. Advances in Neural Information Processing Systems 36 (2024).
- Dettmers et al. (2023) Tim Dettmers, Ruslan A Svirschevski, Vage Egiazarian, Denis Kuznedelev, Elias Frantar, Saleh Ashkboos, Alexander Borzunov, Torsten Hoefler, and Dan Alistarh. 2023. SpQR: A Sparse-Quantized Representation for Near-Lossless LLM Weight Compression. In The Twelfth International Conference on Learning Representations.
- Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv preprint arXiv:1810.04805 (2018).
- Dibia et al. ([n. d.]) Victor Dibia, Adam Fourney, Gagan Bansal, Forough Poursabzi-Sangdeh, Han Liu, and Saleema Amershi. [n. d.]. Aligning Offline Metrics and Human Judgments of Value for Code Generation Models. In Findings of the Association for Computational Linguistics: ACL 2023, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). https://doi.org/10.18653/v1/2023.findings-acl.540
- Ezzini et al. ([n. d.]) Saad Ezzini, Sallam Abualhaija, Chetan Arora, and Mehrdad Sabetzadeh. [n. d.]. Automated handling of anaphoric ambiguity in requirements: a multi-solution study. In Proceedings of the 44th International Conference on Software Engineering. https://doi.org/10.1145/3510003.3510157
- Fatima et al. (2023) Sakina Fatima, Taher A. Ghaleb, and Lionel Briand. 2023. Flakify: A Black-Box, Language Model-Based Predictor for Flaky Tests. IEEE Transactions on Software Engineering 49, 4 (2023), 1912–1927. https://doi.org/10.1109/TSE.2022.3201209
- Feng et al. (2020) Zhangyin Feng, Daya Guo, Duyu Tang, Nan Duan, Xiaocheng Feng, Ming Gong, Linjun Shou, Bing Qin, Ting Liu, Daxin Jiang, and Ming Zhou. 2020. CodeBERT: A Pre-Trained Model for Programming and Natural Languages. In Findings of the Association for Computational Linguistics: EMNLP 2020. Association for Computational Linguistics. https://doi.org/10.18653/v1/2020.findings-emnlp.139
- Fried et al. (2023) Daniel Fried, Armen Aghajanyan, Jessy Lin, Sida Wang, Eric Wallace, Freda Shi, Ruiqi Zhong, Scott Yih, Luke Zettlemoyer, and Mike Lewis. 2023. InCoder: A Generative Model for Code Infilling and Synthesis. In The Eleventh International Conference on Learning Representations. https://openreview.net/forum?id=hQwb-lbM6EL
- Ganesh et al. (2021) Prakhar Ganesh, Yao Chen, Xin Lou, Mohammad Ali Khan, Yin Yang, Hassan Sajjad, Preslav Nakov, Deming Chen, and Marianne Winslett. 2021. Compressing Large-Scale Transformer-Based Models: A Case Study on BERT. Transactions of the Association for Computational Linguistics 9 (2021). https://doi.org/10.1162/tacl_a_00413
- Gou et al. (2021) Jianping Gou, Baosheng Yu, Stephen J Maybank, and Dacheng Tao. 2021. Knowledge distillation: A survey. International Journal of Computer Vision 129, 6 (2021), 1789–1819.
- Guo et al. (2022) Daya Guo, Shuai Lu, Nan Duan, Yanlin Wang, Ming Zhou, and Jian Yin. 2022. UniXcoder: Unified Cross-Modal Pre-training for Code Representation. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics. https://doi.org/10.18653/v1/2022.acl-long.499
- Guo et al. (2021) Daya Guo, Shuo Ren, Shuai Lu, Zhangyin Feng, Duyu Tang, Shujie LIU, Long Zhou, Nan Duan, Alexey Svyatkovskiy, Shengyu Fu, Michele Tufano, Shao Kun Deng, Colin Clement, Dawn Drain, Neel Sundaresan, Jian Yin, Daxin Jiang, and Ming Zhou. 2021. GraphCodeBERT: Pre-training Code Representations with Data Flow. In International Conference on Learning Representations.
- Hendrycks and Gimpel (2016) Dan Hendrycks and Kevin Gimpel. 2016. Gaussian error linear units (gelus). arXiv preprint arXiv:1606.08415 (2016).
- Hindle et al. (2012) Abram Hindle, Earl T. Barr, Zhendong Su, Mark Gabel, and Premkumar Devanbu. 2012. On the naturalness of software. In Proceedings of the 34th International Conference on Software Engineering (ICSE ’12).
- Hinton et al. (2015) Geoffrey E. Hinton, Oriol Vinyals, and Jeffrey Dean. 2015. Distilling the Knowledge in a Neural Network. ArXiv abs/1503.02531 (2015).
- Hou et al. (2023a) Xinyi Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John Grundy, and Haoyu Wang. 2023a. Large language models for software engineering: A systematic literature review. arXiv preprint arXiv:2308.10620 (2023).
- Hou et al. (2023b) Xinying Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John C. Grundy, and Haoyu Wang. 2023b. Large Language Models for Software Engineering: A Systematic Literature Review. ArXiv abs/2308.10620 (2023). https://api.semanticscholar.org/CorpusID:261048648
- Hunger (2005) R. Hunger. 2005. Floating Point Operations in Matrix-vector Calculus. Munich University of Technology, Inst. for Circuit Theory and Signal Processing. https://books.google.ca/books?id=EccIcgAACAAJ
- Husain et al. (2019) Hamel Husain, Ho-Hsiang Wu, Tiferet Gazit, Miltiadis Allamanis, and Marc Brockschmidt. 2019. Codesearchnet challenge: Evaluating the state of semantic code search. arXiv preprint arXiv:1909.09436 (2019).
- Jin et al. (2024) Renren Jin, Jiangcun Du, Wuwei Huang, Wei Liu, Jian Luan, Bin Wang, and Deyi Xiong. 2024. A Comprehensive Evaluation of Quantization Strategies for Large Language Models. arXiv preprint arXiv:2402.16775 (2024).
- Kim et al. (2023) Sehoon Kim, Coleman Hooper, Amir Gholami, Zhen Dong, Xiuyu Li, Sheng Shen, Michael W Mahoney, and Kurt Keutzer. 2023. Squeezellm: Dense-and-sparse quantization. arXiv preprint arXiv:2306.07629 (2023).
- Kim et al. (2022) Sehoon Kim, Sheng Shen, David Thorsley, Amir Gholami, Woosuk Kwon, Joseph Hassoun, and Kurt Keutzer. 2022. Learned Token Pruning for Transformers. In Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’22). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3534678.3539260
- Kobayashi et al. (2020) Goro Kobayashi, Tatsuki Kuribayashi, Sho Yokoi, and Kentaro Inui. 2020. Attention is not only a weight: Analyzing transformers with vector norms. (2020).
- Lacoste et al. (2019) Alexandre Lacoste, Alexandra Luccioni, Victor Schmidt, and Thomas Dandres. 2019. Quantifying the Carbon Emissions of Machine Learning. arXiv preprint arXiv:1910.09700 (2019).
- Lai et al. (2023) Yuhang Lai, Chengxi Li, Yiming Wang, Tianyi Zhang, Ruiqi Zhong, Luke Zettlemoyer, Wen-tau Yih, Daniel Fried, Sida Wang, and Tao Yu. 2023. DS-1000: A natural and reliable benchmark for data science code generation. In International Conference on Machine Learning. PMLR, 18319–18345.
- Le-Cong et al. (2023) Thanh Le-Cong, Duc-Minh Luong, Xuan Bach D. Le, David Lo, Nhat-Hoa Tran, Bui Quang-Huy, and Quyet-Thang Huynh. 2023. Invalidator: Automated Patch Correctness Assessment Via Semantic and Syntactic Reasoning. IEEE Transactions on Software Engineering (2023). https://doi.org/10.1109/TSE.2023.3255177
- Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692 (2019).
- Liu et al. (2021) Zejian Liu, Fanrong Li, Gang Li, and Jian Cheng. 2021. EBERT: Efficient BERT inference with dynamic structured pruning. In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021. 4814–4823.
- Lozhkov et al. (2024) Anton Lozhkov, Raymond Li, Loubna Ben Allal, Federico Cassano, Joel Lamy-Poirier, Nouamane Tazi, Ao Tang, Dmytro Pykhtar, Jiawei Liu, Yuxiang Wei, et al. 2024. StarCoder 2 and The Stack v2: The Next Generation. arXiv preprint arXiv:2402.19173 (2024).
- Lu et al. (2021) Shuai Lu, Daya Guo, Shuo Ren, Junjie Huang, Alexey Svyatkovskiy, Ambrosio Blanco, Colin B. Clement, Dawn Drain, and other. 2021. CodeXGLUE: A Machine Learning Benchmark Dataset for Code Understanding and Generation. CoRR abs/2102.04664 (2021).
- Luccioni et al. (2023) Alexandra Sasha Luccioni, Yacine Jernite, and Emma Strubell. 2023. Power hungry processing: Watts driving the cost of ai deployment? arXiv preprint arXiv:2311.16863 (2023).
- Menzies (2023) Tim Menzies. 2023. Model Review: A PROMISEing Opportunity. In Proceedings of the 19th International Conference on Predictive Models and Data Analytics in Software Engineering (San Francisco, CA, USA) (PROMISE 2023). 64–68. https://doi.org/10.1145/3617555.3617876
- Microsoft Research (2023) Microsoft Research. 2023. CodeXGLUE: General Language Understanding Evaluation benchmark for CODE. https://microsoft.github.io/CodeXGLUE/. Accessed: 2024.
- Min et al. (2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth. 2023. Recent advances in natural language processing via large pre-trained language models: A survey. Comput. Surveys 56, 2 (2023), 1–40.
- Moharil and Sharma (2023) Ambarish Moharil and Arpit Sharma. 2023. TABASCO: A transformer based contextualization toolkit. Science of Computer Programming (2023). https://doi.org/10.1016/j.scico.2023.102994
- Nagel et al. (2021) Markus Nagel, Marios Fournarakis, Rana Ali Amjad, Yelysei Bondarenko, Mart Van Baalen, and Tijmen Blankevoort. 2021. A white paper on neural network quantization. arXiv preprint arXiv:2106.08295 (2021).
- Nijkamp et al. (2023) Erik Nijkamp, Hiroaki Hayashi, Caiming Xiong, Silvio Savarese, and Yingbo Zhou. 2023. Codegen2: Lessons for training llms on programming and natural languages. arXiv preprint arXiv:2305.02309 (2023).
- Niu et al. (2023) Changan Niu, Chuanyi Li, Vincent Ng, Dongxiao Chen, Jidong Ge, and Bin Luo. 2023. An Empirical Comparison of Pre-Trained Models of Source Code (ICSE ’23). https://doi.org/10.1109/ICSE48619.2023.00180
- Paul et al. (2023) Rishov Paul, Md Mohib Hossain, Masum Hasan, and Anindya Iqbal. 2023. Automated program repair based on code review: How do pre-trained transformer models perform? arXiv preprint arXiv:2304.07840 (2023).
- Saad et al. ([n. d.]) Mootez Saad, José Antonio Hernández López, Boqi Chen, Dániel Varró, and Tushar Sharma. [n. d.]. ALPINE replication package. https://github.com/SMART-Dal/FSE25-Alpine-Replication-Package.
- Saad and Sharma (2023) Mootez Saad and Tushar Sharma. 2023. Naturalness of Attention: Revisiting Attention in Code Language Models. (2023).
- Sharma et al. (2024) Tushar Sharma, Maria Kechagia, Stefanos Georgiou, Rohit Tiwari, Indira Vats, Hadi Moazen, and Federica Sarro. 2024. A survey on machine learning techniques applied to source code. Journal of Systems and Software 209 (2024), 111934. https://doi.org/10.1016/j.jss.2023.111934
- Shi et al. (2024) Jieke Shi, Zhou Yang, Hong Jin Kang, Bowen Xu, Junda He, and David Lo. 2024. Greening Large Language Models of Code (ICSE-SEIS’24).
- Shi et al. (2023) Jieke Shi, Zhou Yang, Bowen Xu, Hong Jin Kang, and David Lo. 2023. Compressing Pre-trained Models of Code into 3 MB. In Proceedings of the 37th IEEE/ACM International Conference on Automated Software Engineering (ASE ’22). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3551349.3556964
- Svajlenko et al. (2014) Jeffrey Svajlenko, Judith F Islam, Iman Keivanloo, Chanchal K Roy, and Mohammad Mamun Mia. 2014. Towards a big data curated benchmark of inter-project code clones. In 2014 IEEE International Conference on Software Maintenance and Evolution. IEEE, 476–480.
- Tang et al. (2023) Wei Tang, Mingwei Tang, Minchao Ban, Ziguo Zhao, and Mingjun Feng. 2023. CSGVD: A deep learning approach combining sequence and graph embedding for source code vulnerability detection. J. Syst. Softw. (may 2023). https://doi.org/10.1016/j.jss.2023.111623
- Tian et al. (2023) Haoye Tian, Kui Liu, Yinghua Li, Abdoul Kader Kaboré, Anil Koyuncu, Andrew Habib, Li Li, Junhao Wen, Jacques Klein, and Tegawendé F. Bissyandé. 2023. The Best of Both Worlds: Combining Learned Embeddings with Engineered Features for Accurate Prediction of Correct Patches. ACM Trans. Softw. Eng. Methodol. (2023). https://doi.org/10.1145/3576039
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Ł ukasz Kaiser, and Illia Polosukhin. 2017. Attention is All you Need. In Advances in Neural Information Processing Systems, Vol. 30.
- Wang et al. (2020) Wenhan Wang, Ge Li, Bo Ma, Xin Xia, and Zhi Jin. 2020. Detecting Code Clones with Graph Neural Network and Flow-Augmented Abstract Syntax Tree. In 2020 IEEE 27th International Conference on Software Analysis, Evolution and Reengineering (SANER). IEEE, 261–271.
- Wang et al. (2023b) Xiaorong Wang, Clara Na, Emma Strubell, Sorelle Friedler, and Sasha Luccioni. 2023b. Energy and Carbon Considerations of Fine-Tuning BERT. In Findings of the Association for Computational Linguistics: EMNLP 2023. Association for Computational Linguistics. https://doi.org/10.18653/v1/2023.findings-emnlp.607
- Wang et al. (2023a) Yue Wang, Hung Le, Akhilesh Deepak Gotmare, Nghi DQ Bui, Junnan Li, and Steven CH Hoi. 2023a. CodeT5+: Open code large language models for code understanding and generation. arXiv preprint arXiv:2305.07922 (2023).
- Wang et al. (2021) Yue Wang, Weishi Wang, Shafiq Joty, and Steven C.H. Hoi. 2021. CodeT5: Identifier-aware Unified Pre-trained Encoder-Decoder Models for Code Understanding and Generation. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, 8696–8708. https://doi.org/10.18653/v1/2021.emnlp-main.685
- Xu et al. (2023b) Shengbin Xu, Yuan Yao, Feng Xu, Tianxiao Gu, Jingwei Xu, and Xiaoxing Ma. 2023b. Data Quality Matters: A Case Study of Obsolete Comment Detection. In 2023 IEEE/ACM 45th International Conference on Software Engineering (ICSE). https://doi.org/10.1109/ICSE48619.2023.00074
- Xu et al. (2023a) Yuhui Xu, Lingxi Xie, Xiaotao Gu, Xin Chen, Heng Chang, Hengheng Zhang, Zhengsu Chen, XIAOPENG ZHANG, and Qi Tian. 2023a. QA-LoRA: Quantization-Aware Low-Rank Adaptation of Large Language Models. In The Twelfth International Conference on Learning Representations.
- Yang et al. (2022) Zhou Yang, Jieke Shi, Junda He, and David Lo. 2022. Natural attack for pre-trained models of code. In Proceedings of the 44th International Conference on Software Engineering (ICSE ’22). New York, NY, USA. https://doi.org/10.1145/3510003.3510146
- Yao et al. (2022) Zhewei Yao, Reza Yazdani Aminabadi, Minjia Zhang, Xiaoxia Wu, Conglong Li, and Yuxiong He. 2022. Zeroquant: Efficient and affordable post-training quantization for large-scale transformers. Advances in Neural Information Processing Systems 35 (2022), 27168–27183.
- Zafrir et al. (2019) Ofir Zafrir, Guy Boudoukh, Peter Izsak, and Moshe Wasserblat. 2019. Q8bert: Quantized 8bit bert. In 2019 Fifth Workshop on Energy Efficient Machine Learning and Cognitive Computing-NeurIPS Edition (EMC2-NIPS). IEEE, 36–39.
- Zeng et al. (2022a) Zhengran Zeng, Hanzhuo Tan, Haotian Zhang, Jing Li, Yuqun Zhang, and Lingming Zhang. 2022a. An extensive study on pre-trained models for program understanding and generation. In Proceedings of the 31st ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA 2022). 39–51. https://doi.org/10.1145/3533767.3534390
- Zeng et al. (2022b) Zhengran Zeng, Hanzhuo Tan, Haotian Zhang, Jing Li, Yuqun Zhang, and Lingming Zhang. 2022b. An extensive study on pre-trained models for program understanding and generation (ISSTA 2022). https://doi.org/10.1145/3533767.3534390
- Zhang et al. (2023) Quanjun Zhang, Chunrong Fang, Weisong Sun, Yan Liu, Tieke He, Xiaodong Hao, and Zhenyu Chen. 2023. Boosting Automated Patch Correctness Prediction via Pre-trained Language Model. (2023). https://doi.org/10.48550/ARXIV.2301.12453
- Zhang et al. (2022) Zhaowei Zhang, Hongyu Zhang, Beijun Shen, and Xiaodong Gu. 2022. Diet code is healthy: simplifying programs for pre-trained models of code. In Proceedings of the 30th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE 2022). https://doi.org/10.1145/3540250.3549094
- Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
- Zhou et al. (2019) Yaqin Zhou, Shangqing Liu, Jingkai Siow, Xiaoning Du, and Yang Liu. 2019. Devign: Effective vulnerability identification by learning comprehensive program semantics via graph neural networks. In Advances in Neural Information Processing Systems. 10197–10207.
- Zhu et al. (2023) Xunyu Zhu, Jian Li, Yong Liu, Can Ma, and Weiping Wang. 2023. A survey on model compression for large language models. arXiv preprint arXiv:2308.07633 (2023).